HETAV PATEL

Grade 11,

Find the acceleration due to gravity in a mine of depth 640 m if the value at the surface is 9.800 m s-2. The radius of the earth is 6400 km.

Find the acceleration due to gravity in a mine of depth 640 m if the value at the surface is 9.800 m s-2. The radius of the earth is 6400 km.

Grade:upto college level

1 Answers

Aditi Chauhan
9 years ago
Sol. Let g’ be the acceleration due to gravity in mine. Then g’ = g(1 – d/R) = 9.8 (1 – 640/6400 * 10^3) = 9.8 * 0.9999 = 9.799 m/s^2